a right triangle the sides are related by pythagorean theorem, c^2=a^2+b^2 where c is the hypotenuse . find the hypotenuse when the other 2 sides measurements are 3 feet and 4 feet show work please
Answer by stanbon(75887)

Pythagoras determined (hypotenuse)^2 = (side 1)^2 + (side 2)^2
h^2 = 3^2 + 4^2
h^2 = 9 + 16
h^2 = 25
hypotenuse = 5
--------
Cheers,
Stan H.